Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used iComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ComComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Required 5 ComRequired 1 Required 2 Required 3 Required 4 Required 5 Prepare a schedule of cost of goods manufactured for the year. AssumeRequired 1 Required 2 Required 3 Required 4 Required 5 Compute the unadjusted cost of goods sold for the year. Do not includeRequired 1 Required 2 Required 3 Required 4 Required 5 ---------- Assume that the $38,000 ending balance in Work in Process i

1. Compute the predetermined overhead rate for the year.

2. Compute the amount of underapplied or overapplied overhead for the year.

3. Prepare a schedule of cost of goods manufactured for the year. Assume all raw materials are used in production as direct materials.

4. Compute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answer.

5. Assume that the $38,000 ending balance in Work in Process includes $8,100 of direct materials. Given this assumption, supply the information missing below:
Answer #1
Required 1 :
Predetermined overhead rate = Estimated manufacturing overhead / Estimated direct materials costs = 133500 / 89000 150%
Required 2 :
Beginning raw materials 28000
(+) Purchase of raw materials 132000
Total raw materials available 160000
(-) Ending raw materials 10000
Raw materials used 150000
Overhead applied = Raw materials used * Predetermined overhead rate = 150000 * 150% 225000
Indirect labor 142500
Property taxes 8700
Depreciation of equipment 18000
Maintenance 16000
Insurance 7300
Rent, building 37000
Actual overhead costs 229500
As the overhead applied are less than the actual overhead costs, the overheads are underapplied
Underapplied overhead = Actual overhead costs - Overhead applied = 229500 - 225000 4500
Underapplied Overhead 4500
Required 3 :
Gitano Products
Schedule of cost of goods manufactured
Direct materials :
Beginning raw materials 28000
Purchase of raw materials 132000
   Total raw materials available 160000
Ending raw materials 10000
Raw materials used in production 150000
Direct labor cost 85000
Manufacturing overhead applied 225000
Total manufacturing costs 460000
Beginning work in process 46000
Total cost of work in process 506000
Ending work in process 38000
Cost of goods manufactured 468000
Required 4 :
Cost of goods manufactured 468000
(+) Beginning finished goods 70000
(-) Ending finished goods 60000
Unadjusted cost of goods sold 478000
Required 5 :
Direct materials 8100
Direct labor cost ( 38000 - 8100 - 12150 ) 17750
Manufacturing overhead ( 8100 * 150% ) 12150
Work in process inventory 38000